OpenStudy (anonymous):

how do you find the lcd of 1/(x^2-36),(x+3)/(7x^3-42x^2)

OpenStudy (anonymous):

\[\frac{1}{(x^2-36)}=\frac{1}{(x+6)(x-6)}\] \[\frac{ (x+3) }{ (7x^3-42x^2) }=\frac{ (x+3) }{ 7x^2(x-6) }\]

OpenStudy (anonymous):

what will make the two fractions has the same denominator

OpenStudy (anonymous):

@niho u there?

OpenStudy (anonymous):

sorry..don't really know how this site works

OpenStudy (anonymous):

idk how to do the problem

OpenStudy (anonymous):

what will you multiply the 1st den. with to get a common den. with the 2nd one and what will you muliply the 2nd one with to get a common den

OpenStudy (anonymous):

I really don't know.

OpenStudy (anonymous):

\[\frac{1}{(x+6)(x-6)}*\frac{7x^2}{7x^2}=\frac{7x^2}{7x^2(x+6)(x-6)}\] \[\frac{(x+3)}{7x^2(x -6)}*\frac{(x+6)}{(x+6)}=\frac{(x+3)(x+6)}{7x^2(x+6)(x-6)}\] what is the common den?

OpenStudy (anonymous):

7x^2(x+6)(x-6)

OpenStudy (anonymous):

thats it

OpenStudy (anonymous):

thankyou